The height h in feet of an object after t seconds is given by the function
h=-16t^2+90t+4
how long will it take the object to hit the grown? Round to the nearest thousandth

Answer by stanbon(75887)

The height h in feet of an object after t seconds is given by the function
h=-16t^2+90t+4
how long will it take the object to hit the grown? Round to the nearest thousandth
---------
When it hits the ground its height will be zero.
-------
EQUATION:
-16t^2+90t+4=0
-8t^2+45t+2=0
t=[-45+-sqrt(45^2-4*-8*2)]/-16
t=[-45+-sqrt(2089]/-4
t= 5.669 seconds